The AP is 21,18,15,.............. . Which term of the given A.P. is zero?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
In order to solve this problem assume a term to be zero and then find the number of that term. Get the first term, common difference from the AP given. Complete step-by-step answer: The given AP is 21,18,15,.............. We can clearly see first term is a1= 21 = a ……(1) And common difference can be calculated as a2 - a1=d So from the AP we can say a2 = 18 Therefore, d = 18 – 21 = -3 ……(2) So, d = -3 Let us assume the nth term to be 0. Then, an= 0 ……(3) We know an= a+(n-1)d So we can do, an= 21+(n-1)(-3)=0 ……(From (1), (2) and (3)) 21 + -3n+3=0 24=3n n = 243 n = 8. Hence, the 8th term of the AP is Zero. Note: Whenever you face such types of problems obtain the first term and common difference from the given AP. Here in this question we have assumed the nth term to be zero then we have applied the formula of nth term of the AP. Proceeding like this you will get the right solution.